A steel tank containing one mole of HCl(g) at 500 K is allowed to cool in a Miami, FL warehouse in December where the ambient temperature is 295 K. Calculate the change in the Helmholtz energy ΔA in kJ mol-1, for this process of attaining thermal equilibrium. The molar entropy Sm of HCl(g) over the aforementioned temperature range is given by

Sm / (J K-1 mol-1) = 8.50 × 10-2T + 160

(Reminder: CV,m for HCl(g) is 36.2 J K-1 mol-1 and can be considered constant over the temperature range of analysis.)
Answer #1

Thank you,, happy to help you,, have a very good day,, please like so dr=o As we know mat ; dA= -5dT - POV but cum in condant eq becomes dAs - 5d7 integrate oom sider, we get . 72 + 60T A = -OA = -8.50 810[sco - 295] +2.5015003-295323 . 760€ Sco- 295) OA = -17425 + 13852P7.5 - 32800 GA= 1335062.5 Il mole Tot= 1335.